Circular No. 8839

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 K3 (SIDING SPRING) An apparently asteroidal object discovered by G. Garradd in the course

of the Siding Spring Survey (discovery observation tabulated below) was found by R. H. McNaught on three co-added 0.5-m Uppsala Schmidt telescope images from May 25.35–25.49 UT to show a 6′′ coma that is slightly but certainly diffuse.

2007 UT  α_{2000}  δ_{2000}  Mag.
May 23.43266  11^{h}45^{m}57.^{s}24  −17^{o}47′59′′.4  19.0

The available astrometry (including prediscovery Siding Spring observations by McNaught from Apr. 18), the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-K56.

T = 2008 Apr. 21.735 TT  ω = 23.^{o}561
Ω = 263.299 }2000.0
q = 2.05102 AU  i = 16.295

COMET C/2007 K4 (GIBBS) A. R. Gibbs reports his discovery of a comet on Catalina Sky Survey

images taken with the 0.68-m Schmidt telescope (discovery observation tabulated below), the object noted as having a compact 9′′ coma and a 15′′ tail spanning p.a. 0^{o}–30^{o} on four co-added 30-s unfiltered CCD images taken in good seeing. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, R. H. McNaught writes that his images of this object with the 0.5-m Uppsala Schmidt telescope on May 25.6 UT show a 10′′ coma and a possible tail to the north. Also, R. E. Hill reports that four co-added 30-s Mt. Lemmon images from May 26.35–26.36 show a broad tail ∼ 1′ long in p.a. 10^{o} and a slight coma ∼ 2′′-3′′ in diameter.

2007 UT  α_{2000}  δ_{2000}  Mag.
May 25.32443  16^{h}56^{m}28.^{s}69  −4^{o}37′46′′.0  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-K57.

T = 2007 Nov. 12.567 TT  ω = 211.^{o}884
Ω = 69.590 }2000.0
q = 2.85465 AU  i = 93.350
2007 May 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green